OpenStudy (anonymous):

What is the variance of the following data? If necessary, round your answer to two decimal places. 9, 10, 10, 8, 7, 11, 12, 9 2.25 3.45 5.67 5.89

OpenStudy (anonymous):

First find the expected value: \[1/n (0 + 1 + 2 + 3 + 4.....+n)\] where 0, 1, 2, 3, 4, etc... are the elements of the set and 'n' is the number of elements in the set.

OpenStudy (anonymous):

i dnt understand

OpenStudy (valpey):

First, find the average of each of the elements.\[\frac{9+10+10+...}{8}\]Second, find the average of the squares of each of the elements. \[\frac{9^2+10^2+10^2+...}{8}\]The 8 on the bottom is because there are 8 elements.

OpenStudy (anonymous):

so i do the number squared and just add them all like that then divide by 8

OpenStudy (valpey):

Yeah, that is how you get to the average of the squares. Then, to get the variance you take: (the average of the squares) - the square of the average

OpenStudy (anonymous):

okaii thank yu is that all i need too know

OpenStudy (valpey):

So if my numbers were 1, 2, 3, 4 the average would be 2.5. The average of the squares would be (1+4+9+16)/4 = 7.5 and the variance would be 7.5 - (2.5)^2 = 1.25

OpenStudy (anonymous):

i tried doing what you said the answer didnt come out right